### v1

@@
[…]
Hu et al.
[…]
CLO with Bandit Feedback
[…]
Contextual Linear Optimization with Bandit Feedback
[…]
Contextual linear optimization (CLO) uses predictive observations to reduce uncertainty in random cost coefficients and thereby improve average-cost performance. An example is stochastic shortest path with random edge costs (e.g., traffic) and predictive features (e.g., lagged traffic, weather). Existing work on CLO assumes the data has fully observed cost coefficient vectors, but in many applications we can only see the realized cost of a historical decision, that is, just one projection of the random cost coefficient vector, to which we refer as bandit feedback. We study a class of algorithms for CLO with bandit feedback, which we term induced empirical risk minimization (IERM), where we fit a predictive model to directly optimize downstream performance of the policy it induces. We show a fast-rate regret bound for IERM that allows for misspecified model classes and flexible choices of the optimization estimate, and we develop computationally tractable surrogate losses. A byproduct of our theory of independent interest is fast-rate regret bound for IERM with full feedback and misspecified policy class. We compare the performance of different modeling choices numerically using a stochastic shortest path example and provide practical insights from the empirical results.
@@ 1 Introduction
CLO has been the focus of much recent work (e.g., El Balghiti et al. 2019, Elmachtoub and Grigas 2022, Hu et al. 2022, Sadana et al. 2024, Mandi et al. 2023), due to its relevance to data-driven operational decision making in many applications such as network optimization, portfolio optimization, product recommendation, etc. The key question in these is how to use data to learn a good policy, $\pi:\mathbb{R}^{p}\to\mathcal{Z}$, mapping feature observations to effective decisions. In all of these literature, the data considered are draws $(X_{i},Y_{i})$ from the joint distribution of $(X,Y)$. Crucially, this assumes that we fully observe the random cost vector $Y_{i}$’s, meaning that we fully know the corresponding costs $Y_{i}^{\top}z$ for any potential decision $z$. This may be unrealistic in many applications.
@@
Consider for example the instantiation of Eq. 1 as a stochastic shortest path problem of transporting one unit from a start node to a desired end node, where $Y$ contains the travel time on each of $d$ edges, $z$ the binary decision whether to traverse each edge, and $\mathcal{Z}$ constrains flow conservation. The data we would need in order to apply existing approaches here would consist of simultaneous observations of the travel times on every single edge, encapsulated in the elements of $Y_{i}$’s. However, such ideal observations may not be available. Indeed, modern travel time prediction models are often based on data from historical individual trips. For example, such is the case with Uber Movement data (Sun et al. 2020, Maass et al. 2020), which is based on the total length of historical rides. Namely, instead of observing the entire $Y_{i}$, we only observe the total travelling time $C_{i}=Y_{i}^{\top}Z_{i}$ for the path $Z_{i}$ in a historical trip $i$. We term this observation model bandit feedback as it corresponds to observing the cost only of a given decision and not the counterfactual cost of any alternative decision, as in bandit problems (Lattimore and Szepesvári 2020).
@@
In this paper, we study the contextual linear optimization problem with bandit feedback and make several contributions. First, we adapt the end-to-end induced empirical risk minimization approach to the bandit feedback setting. This approach allows us to optimize decision policies by directly targeting the decision objective. We provide three different methods to identify the expected cost of any given policy and show how to estimate the expected decision cost from bandit-feedback data. Second, we derive upper bounds on the regret (i.e., decision sub-optimality) of the policy that minimizes the estimated objective. Our regret analysis accounts for the misspecification of the policy class and incorporates a margin condition that potentially enables a faster regret rate. This significantly extends existing theory for full-feedback CLO, and as a byproduct, we provide a novel fast rate bound for full-feedback CLO with misspecification. Finally, we demonstrate that existing surrogate losses for full-feedback CLO (such as the well-known SPO+ loss in Elmachtoub and Grigas (2022)) can be adapted to our setting, enabling efficient policy optimization. We empirically test this in simulated shortest path problems and provide practical insights.
@@ 1.1 Background: The Full Feedback Setting
In contrast, some recent literature propose to integrate the estimation and the optimization, directly searching for a decision policy to minimize the decision cost. Following Hu et al. (2022), we consider an IERM formulation that minimizes the sample average cost over the class of plug-in policies induced by $\mathcal{F}$:
[…]
This approach is end-to-end in that that it directly targets the decision-making objective. Recent literature demonstrate that end-to-end approaches like IERM often outperform the ETO approach in terms of the decision quality, in particular for a misspecified $\mathcal{F}$ class that fails to include the truth $f_{0}$ (e.g., Elmachtoub and Grigas 2022, Hu et al. 2022, Elmachtoub et al. 2023). While Eq. 3 involves a challenging bi-level optimization due to the definition of $\Pi_{\mathcal{F}}$, practical computational approximations have been proposed (Elmachtoub and Grigas 2022, Huang and Gupta 2024).
@@ 1.2 Our Problem: CLO with Bandit Feedback
We now formally define the data generating process in the bandit feedback setting. Assume we have $n$ data points $\mathcal{D}=\left\{\left(X_{1},Z_{1},C_{1}\right),\dots,\left(X_{n},Z_{n},C_{n}\right)\right\}$ that are independent draws from a distribution $P$ on $(X,Z,C)$ generated in the following way. We first draw $(X,Z,Y)$ where the $(X,Y)$ distribution is as in the full feedback setting. Then we set $C=Z^{\top}Y$ and omit the full $Y$. Below we impose some basic assumptions on the generating process of the historical decision $Z$.
{assumption}
@@
The ignorability condition is a common assumption that plays a fundamental role in learning with partial feedback and causal inference (see the literature in Section 1.3). This condition requires that the assignments of the historical decisions $Z$ do not depend on any unobserved factor potentially dependent on $Y$. The coverage condition requires that given the covariates $X$, historical decisions $Z$ can explore all directions of the linear span of the constraint set $\mathcal{Z}$; otherwise it may be impossible to evaluate certain policies from the observed costs of historical decisions. This is an analogue of the overlap assumption in learning with partial feedback and causal inference Imbens and Rubin (2015).
@@ 1.3 Background: Off-Policy Evaluation and Optimization
An important special case is when $\mathcal{Z}^{\angle}=\{(1,0,\dots,0),\dots,(0,\dots,0,1)\}$ is the canonical basis and $\mathcal{Z}$ is the simplex. This corresponds to choosing one of $d$ actions (or randomizing between them). The full-feedback problem in this case is cost-sensitive classification (Elkan 2001). In the bandit feedback case, with the restriction $Z\in\mathcal{Z}^{\angle}$, it is known as a logged or offline contextual bandit, and a long line of literature studies the problems of estimating and optimizing $V(\pi)$ Dudík et al. (2014), Kallus (2017, 2018), Athey and Wager (2021), Zhou et al. (2023), Kitagawa and Tetenov (2018), Swaminathan and Joachims (2015a, b, c), Zhao et al. (2012), Zhou et al. (2017), Zhang et al. (2012), Kallus (2021), Kallus et al. (2022), Si et al. (2023). It is closely related to causal inference, viewing each component of $Y$ as the potential outcome for the corresponding action (or treatment), where we only see the realized outcome $C$ corresponding to the factual action with index $1$ in $Z\in\mathcal{Z}^{\angle}$ and see nothing about counterfactual actions $\mathcal{Z}^{\angle}\backslash\{Z\}$. Two key differentiating characteristics of our problem with a general polytope constraint $\mathcal{Z}$ is the opportunity to leverage the linear cost structure to extrapolate from the costs of historical decisions to costs of other decisions, and the presence of non-trivial constraints on the output of decision policies.
Going beyond just finite discrete arms, Kallus and Zhou (2018), Kallus and Uehara (2020), Sachdeva et al. (2024) consider the logged contextual bandit with a continuum of arms, which is not generally representable in terms of the problem in Section 1.2. These works make no restrictions on the relationship between the expected potential outcome and the corresponding action except generic smoothness, and leverage nonparametric approaches such as kernel weighting. Closest to our work is Chernozhukov et al. (2019), which imposes a semiparametric assumption on this relationship. This includes our problem (Section 1.2) as a special case under the restriction that, conditional on $X$, expected outcomes are linear in actions (note that our $\mathbb{E}\left[C\mid Z,X\right]$ is linear in $Z$ according to ignorability in Section 1.2). Relative to this work, our unique contributions are the use of induced policy classes to naturally deal with the constraints in $\mathcal{Z}$, the adaptation of computational approximation such as SPO+, and obtaining the fast rates for the regret under misspecification.
@@ 2 Induced Empirical Risk Minimization with Bandit Feedback
Define $\Sigma_{0}(X)=\mathbb{E}\left[ZZ^{\top}\mid X\right]$ and recall $f_{0}(X)=\mathbb{E}\left[Y\mid X\right]$, which we call as nuisance functions as they will be used to identify the policy value. Let $\theta(x,z,c;f,\Sigma)$ be a score function such that
@@
where $\mathbb{E}_{P}$ denotes taking expectation over $P$. Proposition 2.1 summarizes a few possible choices.
Proposition 2.1
@@
The doubly robust identification in Proposition 2.1 is a generalization of the identification in Chernozhukov et al. (2019) in that we allow for rank-deficient $\Sigma_{0}$ by using a pseudo-inverse. Otherwise, our work significantly differs from Chernozhukov et al. (2019) in terms of policy class specification, computation and theoretical analyses (see Section 1.3). Note that identification in Proposition 2.1 exploits the linearity of the decision cost, which significantly improve upon algorithms that ignore the linear structure and naively extend existing offline bandit learning using discrete actions ${\mathcal{Z}^{\angle}}$; see Sections 5 and 8.2 for numerical evidence.
It remains to estimate Eq. 5, since we know neither $P$ nor the nuisance functions $f_{0},\Sigma_{0}$. Following the approach of Athey and Wager (2021), Zhou et al. (2023) for logged contextual bandits, we adapt the cross-fitting procedure advocated in Chernozhukov et al. (2018). For simplicity, we focus on the twofold version and assume $n$ is even. The extension to $K$-fold version is straightforward.
@@
Remark 2.3 (Estimation of $\Sigma_{0}$)
There are multiple ways to estimate $\Sigma_{0}$. For example, Chernozhukov et al. (2019) suggest estimating $\Sigma_{0}$ by running a multi-task regression for all $(j,k)$ entries to the matrix over some appropriate hypothesis spaces $\mathcal{S}_{jk}$: $\hat{\Sigma}=\arg\min_{\Sigma_{11}\in\mathcal{S}_{11},\dots,\Sigma_{dd}\in\mathcal{S}_{dd}}\sum_{i\in\mathcal{D}}\|Z_{i}Z_{i}^{\top}-\Sigma(X_{i})\|_{Fro}^{2}$. To ensure a positive semi-definite estimator, we may posit each hypothesis $\Sigma$ to be the outer product of some matrix-valued hypothesis of appropriate dimension. Alternatively, we may only need to consider finitely many feasible decisions $z_{1},\dots,z_{m}$, such as the feasible paths for stochastic shortest path problems (see experiments in Section 5) or more generally the vertices in $\mathcal{Z}^{\angle}$. Then we can first estimate the propensity scores $e_{0}(z\mid X)$ for $z=z_{1},\dots,z_{m}$ using an suitable estimator $\hat{e}(z\mid X)$ and then estimate $\Sigma_{0}(X)=\sum_{j=1}^{m}z_{j}z_{j}^{\top}e_{0}(z_{j}\mid X)$ by $\hat{\Sigma}(X)=\sum_{j=1}^{m}z_{j}z_{j}^{\top}\hat{e}(z_{j}\mid X)$.
@@ 3 Theoretical Analysis
Throughout Section 3, we impose two assumptions. Section 3 concerns the algorithms we use to estimate the nuisance functions $f_{0},\Sigma_{0}$. It requires that the nuisance estimates are close enough to their true values, and $\|{\theta(x,z,c;\hat{f},\hat{\Sigma})}\|$ is bounded. {assumption}[Nuisance Estimation] The nuisance estimators $\hat{f},\hat{\Sigma}$ trained on a sample of size $n$ satisfy that $\|\theta(x,z,c;\hat{f},\hat{\Sigma})\|\leq\Theta$ for all $x,z,c$, and for any $\delta\in(0,1)$ and $\pi\in\Pi_{\mathcal{F}}$, with probability at least $1-\delta$,
@@
Section 3, which we term the margin condition, controls the density of the sub-optimality gap near zero in the CLO problem instance and allows us to get faster regret rates. This type of condition was originally considered in the binary classification literature (Tsybakov 2004, Audibert and Tsybakov 2007). It is recently extended to contextual linear optimization by Hu et al. (2022), which we describe below. {assumption}[Margin Condition] Let $\mathcal{Z}^{*}(x)=\arg\min_{z\in\mathcal{Z}}f_{0}(x)^{\top}z$, and $\Delta(x)=\inf_{z\in\mathcal{Z}^{\angle}\backslash\mathcal{Z}^{*}(x)}f_{0}(x)^{\top}z-\inf_{z\in\mathcal{Z}}f_{0}(x)^{\top}z$ if $\mathcal{Z}^{*}(x)\neq\mathcal{Z}$ and $\Delta(x)=0$ otherwise. Assume for some $\alpha,\gamma\geq 0$,
@@
Lemmas 4 and 5 in Hu et al. (2024) show that Section 3 generally holds with $\alpha=1$ for sufficiently well-behaved $f_{0}$ and continuous $X$ and with $\alpha=\infty$ for discrete $X$. Moreover, any CLO instance trivially satisfies $\alpha=0$. Overall, a larger $\alpha$ means that the sub-optimality gap between the best and second-best decisions tends to be large in more contexts, so it is easier to distinguish the optimal decisions from others. We will show that a larger $\alpha$ parameter could lead to faster regret rates.

### v2

@@
Abstract
Contextual linear optimization (CLO) uses predictive contextual features to reduce uncertainty in random cost coefficients and thereby improve average-cost performance. An example is the stochastic shortest path problem with random edge costs (e.g., traffic) and contextual features (e.g., lagged traffic, weather). Existing work on CLO assumes the data has fully observed cost coefficient vectors, but in many applications, we can only see the realized cost of a historical decision, that is, just one projection of the random cost coefficient vector, to which we refer as bandit feedback. We study a class of offline learning algorithms for CLO with bandit feedback, which we term induced empirical risk minimization (IERM), where we fit a predictive model to directly optimize the downstream performance of the policy it induces. We show a fast-rate regret bound for IERM that allows for misspecified model classes and flexible choices of the optimization estimate, and we develop computationally tractable surrogate losses. A byproduct of our theory of independent interest is fast-rate regret bound for IERM with full feedback and misspecified policy class. We compare the performance of different modeling choices numerically using a stochastic shortest path example and provide practical insights from the empirical results.
@@ 1 Introduction
CLO has been the focus of much recent work [e.g., 9, 11, 14, 30, 27], due to its relevance to data-driven operational decision making in many applications such as network optimization, portfolio optimization, product recommendation, etc. The key question in these is how to use data to learn a good policy, $\pi:\mathbb{R}^{p}\to\mathcal{Z}$, mapping feature observations to effective decisions. All of this literature studies an offline setting, where a batch of existing data $(X_{i},Y_{i})$ for $i=1,\dots,n$ is observed. These data are viewed as random draws from the joint distribution of $(X,Y)$, and the goal is to learn an effective decision policy from them. Crucially, this assumes that we fully observe the random cost vector $Y_{i}$’s, meaning that we fully know the corresponding costs $Y_{i}^{\top}z$ for any potential decision $z$. This may be unrealistic in many applications.
@@
Consider for example the instantiation of Eq. 1 as a stochastic shortest path problem of transporting one unit from a start node to a desired end node, where $Y$ contains the travel time on each of $d$ edges, $z$ the binary decision whether to traverse each edge, and $\mathcal{Z}$ constrains flow conservation. The data we would need in order to apply existing approaches here would consist of simultaneous observations of the travel times on every single edge, encapsulated in the elements of $Y_{i}$’s. However, such ideal observations may not be available. Indeed, modern travel time prediction models are often based on data from historical individual trips. For example, such is the case with Uber Movement data [33, 26], which is based on the total length of historical rides. Namely, instead of observing the entire $Y_{i}$, we only observe the total travelling time $C_{i}=Y_{i}^{\top}Z_{i}$ for the path $Z_{i}$ in a historical trip $i$. We term this observation model bandit feedback as it corresponds to observing the cost only of a given decision and not the counterfactual cost of any alternative decision, as in bandit problems [25].
@@
In this paper, we study the contextual linear optimization problem with bandit feedback and make several contributions. First, we adapt the end-to-end induced empirical risk minimization approach to the bandit feedback setting. This approach allows us to optimize decision policies by directly targeting the decision objective. We provide three different methods to identify the expected cost of any given policy and show how to estimate the expected decision cost from bandit-feedback data. Second, we derive upper bounds on the regret (i.e., decision sub-optimality) of the policy that minimizes the estimated objective. Our regret analysis accounts for the misspecification of the policy class and incorporates a margin condition that potentially enables a faster regret rate. This significantly extends existing theory for full-feedback CLO, and as a byproduct, we provide a novel fast rate bound for full-feedback CLO with misspecification. Finally, we demonstrate that existing surrogate losses for full-feedback CLO (such as the well-known SPO+ loss in [11]) can be adapted to our setting, enabling efficient policy optimization. We empirically test this in simulated shortest path problems and provide practical insights.
@@ 1.1 Background: The Full Feedback Setting
In contrast, some recent literature propose to integrate the estimation and the optimization, directly searching for a decision policy to minimize the decision cost. Following [14], we consider an IERM formulation that minimizes the sample average cost over the class of plug-in policies induced by $\mathcal{F}$:
$\textstyle\hat{\pi}^{\text{F}}\in\arg\min_{\pi\in\Pi_{\mathcal{F}}}\frac{1}{n}\sum_{i=1}^{n}Y_{i}^{\top}\pi(X_{i}),~{}~{}\text{where }\Pi_{\mathcal{F}}=\left\{\pi_{f}:f\in\mathcal{F}\right\}.$ (3)
This approach is end-to-end in that it directly targets the decision-making objective. Recent literature demonstrates that end-to-end approaches like IERM often outperform the ETO approach [12]. The benefits of end-to-end approaches are significant especially in the misspecified setting - that is, when the function class $\mathcal{F}$ fails to contain the true expected cost function $f_{0}$, and the policy class $\Pi_{\mathcal{F}}$ does not include the globally optimal decision policy $\pi_{f_{0}}$ [e.g., 11, 14]. While Eq. 3 involves a challenging bi-level optimization due to the definition of $\Pi_{\mathcal{F}}$, practical computational approximations have been proposed [11, 16].
@@ 1.2 Our Problem: CLO with Bandit Feedback
We now formally define the data generating process in the bandit feedback setting. Assume we have an offline dataset consisting of $n$ data points $\mathcal{D}=\left\{\left(X_{1},Z_{1},C_{1}\right),\dots,\left(X_{n},Z_{n},C_{n}\right)\right\}$ that are independent draws from a distribution $P$ on $(X,Z,C)$ generated in the following way. We first draw $(X,Z,Y)$ where the $(X,Y)$ distribution is as in the full feedback setting and $Z$ is generated according to some historical decision policies for data collection. Then we set $C=Z^{\top}Y$ and omit the full $Y$. Below we impose some basic assumptions on the generating process of the historical decision $Z$.
Assumption 1.
@@
The ignorability condition is a common assumption that plays a fundamental role in learning with partial feedback and causal inference (see the literature in Section 1.3). This condition requires that the assignments of the historical decisions $Z$ do not depend on any unobserved factor potentially dependent on $Y$. The coverage condition requires that given the covariates $X$, historical decisions $Z$ can explore all directions of the linear span of the constraint set $\mathcal{Z}$; otherwise it may be impossible to evaluate certain policies from the observed costs of historical decisions. This is an analogue of the overlap assumption in learning with partial feedback and causal inference [17].
@@ 1.3 Background: Off-Policy Evaluation and Optimization
An important special case is when $\mathcal{Z}^{\angle}=\{(1,0,\dots,0),\dots,(0,\dots,0,1)\}$ is the canonical basis and $\mathcal{Z}$ is the simplex. This corresponds to choosing one of $d$ actions (or randomizing between them). The full-feedback problem in this case is cost-sensitive classification [10]. In the bandit feedback case, with the restriction $Z\in\mathcal{Z}^{\angle}$, it is known as a logged or offline contextual bandit, and a long line of literature studies the problems of estimating and optimizing $V(\pi)$ [8, 18, 19, 1, 45, 24, 34, 35, 36, 43, 44, 42, 20, 23, 32]. It is closely related to causal inference, viewing each component of $Y$ as the potential outcome for the corresponding action (or treatment), where we only see the realized outcome $C$ corresponding to the factual action with index $1$ in $Z\in\mathcal{Z}^{\angle}$ and see nothing about counterfactual actions $\mathcal{Z}^{\angle}\backslash\{Z\}$. Two key differentiating characteristics of our problem with a general polytope constraint $\mathcal{Z}$ are the opportunity to leverage the linear cost structure to extrapolate from the costs of historical decisions to costs of other decisions, and the presence of non-trivial constraints on the output of decision policies.
Going beyond just finite discrete arms, [22, 21, 29] consider the logged contextual bandit with a continuum of arms, which is not generally representable in terms of the problem in Section 1.2. These works make no restrictions on the relationship between the expected potential outcome and the corresponding action except generic smoothness, and leverage nonparametric approaches such as kernel weighting. Closest to our work is [6], which imposes a semiparametric assumption on this relationship. This includes our problem (Section 1.2) as a special case under the restriction that, conditional on $X$, expected outcomes are linear in actions (note that our $\mathbb{E}\left[C\mid Z,X\right]$ is linear in $Z$ according to ignorability in Section 1.2). Relative to this work, our unique contributions are the use of induced policy classes to naturally deal with the constraints in $\mathcal{Z}$, the adaptation of computational approximation such as SPO+, and obtaining the fast rates for the regret under misspecification.
@@ 2 Induced Empirical Risk Minimization with Bandit Feedback
Define $\Sigma_{0}(X)=\mathbb{E}\left[ZZ^{\top}\mid X\right]$ and recall $f_{0}(X)=\mathbb{E}\left[Y\mid X\right]$ , which we will use to characterize the policy value $V(\pi)$. We refer to these functions as nuisance functions following the existing literature on off-policy evaluation and learning in contextual bandits or reinforcement learning [39], because they are not directly used for decision-making but serve merely as intermediaries for evaluating the policy value. Let $\theta(x,z,c;f,\Sigma)$ be a score function such that
@@
where $\mathbb{E}_{P}$ denotes taking expectation over $P$. Proposition 1 summarizes a few possible choices.
Proposition 1.
@@
The doubly robust identification in Proposition 1 is a generalization of the identification in [6] in that we allow for rank-deficient $\Sigma_{0}$ by using a pseudo-inverse. Otherwise, our work significantly differs from [6] in terms of policy class specification, computation and theoretical analyses (see Section 1.3). Note that identification in Proposition 1 exploits the linearity of the decision cost, which significantly improve upon algorithms that ignore the linear structure and naively extend existing offline bandit learning using discrete actions ${\mathcal{Z}^{\angle}}$; see Sections 5 and C.2 for numerical evidence.
It remains to estimate Eq. 5, since we know neither $P$ nor the nuisance functions $f_{0},\Sigma_{0}$. Following the approach of [1, 45] for logged contextual bandits, we adapt the cross-fitting procedure advocated in [5]. For simplicity, we focus on the twofold version and assume $n$ is even. The extension to $K$-fold version is straightforward.
@@
Remark 2 (Estimation of $\Sigma_{0}$).
There are multiple ways to estimate $\Sigma_{0}$. For example, [6] suggest estimating $\Sigma_{0}$ by running a multi-task regression for all $(j,k)$ entries to the matrix over some appropriate hypothesis spaces $\mathcal{S}_{jk}$: $\hat{\Sigma}=\arg\min_{\Sigma_{11}\in\mathcal{S}_{11},\dots,\Sigma_{dd}\in\mathcal{S}_{dd}}\sum_{i\in\mathcal{D}}\|Z_{i}Z_{i}^{\top}-\Sigma(X_{i})\|_{Fro}^{2}$. To ensure a positive semi-definite estimator, we may posit each hypothesis $\Sigma$ to be the outer product of some matrix-valued hypothesis of appropriate dimension. Alternatively, we may only need to consider finitely many feasible decisions $z_{1},\dots,z_{m}$, such as the feasible paths for stochastic shortest path problems (see experiments in Section 5) or more generally the vertices in $\mathcal{Z}^{\angle}$. Then we can first estimate the propensity scores $e_{0}(z\mid X)$ for $z=z_{1},\dots,z_{m}$ using an suitable estimator $\hat{e}(z\mid X)$ and then estimate $\Sigma_{0}(X)=\sum_{j=1}^{m}z_{j}z_{j}^{\top}e_{0}(z_{j}\mid X)$ by $\hat{\Sigma}(X)=\sum_{j=1}^{m}z_{j}z_{j}^{\top}\hat{e}(z_{j}\mid X)$.
@@ 3 Theoretical Analysis
Throughout Section 3, we impose two assumptions. Assumption 2 concerns the algorithms we use to estimate the nuisance functions $f_{0},\Sigma_{0}$. It requires that the nuisance estimates are close enough to their true values, and $\|{\theta(x,z,c;\hat{f},\hat{\Sigma})}\|$ is bounded.
Assumption 2 (Nuisance Estimation).
The nuisance estimators $\hat{f},\hat{\Sigma}$ trained on a sample of size $n$ satisfy that $\|\theta(x,z,c;\hat{f},\hat{\Sigma})\|\leq\Theta$ for all $x,z,c$, and for any $\delta\in(0,1)$ and $\pi\in\Pi_{\mathcal{F}}$, with probability at least $1-\delta$,
@@
Assumption 3, which we term the margin condition, controls the density of the sub-optimality gap near zero in the CLO problem instance and allows us to get faster regret rates. This type of condition was originally considered in the binary classification literature [38, 2]. It is recently extended to contextual linear optimization by [14], which we describe below.
Assumption 3 (Margin Condition).
Let $\mathcal{Z}^{*}(x)=\arg\min_{z\in\mathcal{Z}}f_{0}(x)^{\top}z$, and $\Delta(x)=\inf_{z\in\mathcal{Z}^{\angle}\backslash\mathcal{Z}^{*}(x)}f_{0}(x)^{\top}z-\inf_{z\in\mathcal{Z}}f_{0}(x)^{\top}z$ if $\mathcal{Z}^{*}(x)\neq\mathcal{Z}$ and $\Delta(x)=0$ otherwise. Assume for some $\alpha,\gamma\geq 0$,
@@
Lemmas 4 and 5 in [15] show that Assumption 3 generally holds with $\alpha=1$ for sufficiently well-behaved $f_{0}$ and continuous $X$ and with $\alpha=\infty$ for discrete $X$. Moreover, any CLO instance trivially satisfies $\alpha=0$. Overall, a larger $\alpha$ means that the sub-optimality gap between the best and second-best decisions tends to be large in more contexts, so it is easier to distinguish the optimal decisions from others. We will show that a larger $\alpha$ parameter could lead to faster regret rates.